OpenStudy (schoolgeek15):

The sum of two numbers is 52. The greater number is 4 more than the smaller number. Which equation can be used to solve for the smaller number? x − (x + 4) = 52 x + (x + 4) = 52 x(x + 4) = 52 x(x − 4) = 52 think its B can someone check my answer?

OpenStudy (danjs):

yeah i agree

OpenStudy (schoolgeek15):

Thanks :)

OpenStudy (danjs):

X+Y=52 and Y=X+4

OpenStudy (danjs):

X + (X+4)=52 X=24 Y=28

OpenStudy (schoolgeek15):

So its B?

OpenStudy (danjs):

right

OpenStudy (schoolgeek15):

thx